# Clarifying the effect of mean subtraction on Dynamic Mode Decomposition

### v5

@@ 5.1.2 Addressing a persistent fragility with ensemble studies
To paraphrase, $\mathbf{Z}$ corresponds to sampling $\boldsymbol{\mathbf{\psi}}$ on the first $\theta+d+1$ elements of the trajectory. Subsequently, we form $\mathbf{Z}_{\rm d-delayed}$, perform DMD [^11] A relative threshold of $10^{-8}$ is used in all pseudo-inverse computations. and compute Relative distance to DFT. This calculation is, then, iterated over every pair $(\theta,d)$ in the parameter regime of interest. Finally, we also repeat this parametric sweep, at a higher level, over many trajectories [^12] We also switch the dictionary for each trajectory for an added level of randomness.. Consequently, we can study the statistics of Relative distance to DFT for every pair of Companion-order ($\theta$) and #$\#[delays]~{}(d)$ being tested. This is facilitated by box-whisker plots of this indicator with respect to Companion-order.
We test the major results on a variety of dynamical systems of increasing complexity. We primarily work with linear-time invariant (LTI) systems since their Koopman spectra are easily determined and the conditions required in Corollary 5.1 can be met. We also validate our theory using the Van der Pol oscillator and the lid-driven cavity from [2]. In contrast to the LTI systems, we have no guarantee of meeting the requisites in Corollary 5.1 nor do we possess complete knowledge of the associated Koopman spectra. Nonetheless, we can still use the contrapositive of Corollary 5.1 to get a lower bound on the complexity of the underlying Koopman mode expansion.
@@
$\{\,\lambda_{i}\,\}_{i=1}^{r}$$\{z\,|\,z^{7}=1\}$ $|\lambda_{i}|=1,~{}\forall\,i$ $|\lambda_{i}|=1,~{}i~{}=~{}1,2,3,4$
$|\lambda_{5}|=0.97,~{}|\lambda_{6}|=0.93,~{}|\lambda_{7}|=0.87$
@@
For the unitary systems $\textrm{LTI}_{1a}$ and $\textrm{LTI}_{1b}$, the expected trend is clearly visible in sub-figures (a) and (b) of Fig. 1. Relative distance to DFT is orders of magnitude larger when $\theta$ crosses the critical value of $7$. Moreover, the shadowing of this feature by the pertinent inter-quartile ranges (IQRs), which are denoted by boxes colored according to the number of delays taken, suggests that the observation is independent of the choice of initial condition.
@@
The computations with $\textrm{LTI}_{3}$ are not in exact agreement with the theoretical claims, particularly when $\theta$ matches $r$. Recall that Corollary 5.1 predicts the same trend for $\textrm{LTI}_{3}$ and $\textrm{LTI}_{1b}$ since neither has a Koopman eigenvalue at $1$. In sub-figure (c) of Fig. 1, we observe that DMD-DFT equivalence appears to hold for $\theta$ less than 7 and is absent when $\theta$ exceeds 7. However, when $\theta$ is 7, we see non-equivalence of mean-subtracted DMD and DFT, in contrary to the theoretical prediction of equivalence. We believe that this is a numerical issue arising from the three Koopman eigenvalues of this system that lie inside the unit disc (see Table 2). These eigenvalues degrade the conditioning of $\mathbf{Z}_{\rm d-delayed}$ [^13] As compared to its realizations for $\textrm{LTI}_{1a}$ and $\textrm{LTI}_{1b}$ in Algorithm 3 and, thus, perturb Relative distance to DFT.
@@
Figure 1: For the linear time-invariant (LTI) systems described by Eq. 19, box plots of Relative distance to DFT reveal its dependence on the model order $(\theta)$ and the number of time delays (Colour-coded). The magenta line indicates the system order $r$, which is $7$ for all three examples. When at least 6 time delays are taken, $\textrm{LTI}_{1a}(\textrm{LTI}_{1b})$ displays a plateau starting at a Companion-order of 7 (8) as predicted by Corollary 5.1. Although this plateau begins earlier than expected ( at $\theta=7$ instead of $\theta=8$) for $\textrm{LTI}_{3}$, the overall trend conforms to the theoretical prediction of Corollary 5.1.
@@
When the assumptions (3.8) and (3.10) are not known to hold, we can still use the contrapositive of Corollary 5.1 to provide a lower bound on the number of Koopman modes constituting our dictionary $\boldsymbol{\mathbf{\psi}}$. First, we note that if the predictions of Corollary 5.1 are inconsistent with a numerical study that has taken sufficient time delays, then, at-least one of the assumptions (3.8) and (3.10) must be invalid. In practice, (3.10) holds true because we can restrict our attention to only those KEFs whose zero level set does not contain $\boldsymbol{\mathbf{\upsilon}}_{1}$. This change of perspective is permissible because the reduced collection of KEFs is equally capable of generating the training set, $\mathbf{Z}$, through time-sequential observations. Since (3.8) can be interpreted as requiring $\boldsymbol{\mathbf{\psi}}$ possess only $r$ Koopman modes, its negation, over the specific guesses of $r$ described below, can give a lower bound on the number of Koopman modes comprising $\boldsymbol{\mathbf{\psi}}$. Thus, when the delay embedding dimension, $d$, is large enough, numerical observations of Relative distance to DFT that are inconsistent with Corollary 5.1 can inform the number of Koopman modes necessary to represent $\boldsymbol{\mathbf{\psi}}$.
The contrapositive of Corollary 5.1 can be computationally leveraged by assuming an upper-bound on the number of Koopman modes, performing the ensemble experiment described in Section 5.1.2 and using errors, if any, in the predictions of Corollary 5.1 to negate the presumed upper-bound. In each forthcoming numerical experiment, we begin with the assumption that our observables are non-redundantly spanned by $r$ distinct KEFs. Although $r$ is unknown, we assume that it is finite and has a known upper bound of $r_{\rm max}$. This may be a theoretically informed cap or, if little is known about the system, simply the maximum model order that is palatable from a modeling perspective. Suppose we take at least $r_{\rm max}-1$ time delays i.e., $d\geq r_{\max}-1$. Then, by Corollary 5.1, Relative distance to DFT must behave like a step function with respect to the model order ($\theta$). It must be near zero for model orders less than $r+1$ ($r$ if 1 is a Koopman eigenvalue) and significantly higher for larger model orders. The absence of a step-like behavior for $\theta$ less than or equal to $r_{\max}+1$ would, then, mean that at-least $r_{\rm max}+1$ distinct KEFs are required to represent the span of our observables i.e., $\boldsymbol{\mathbf{\psi}}$ has at-least $r_{\max}+1$ Koopman modes. In contrast, the presence of a step-like behavior, in agreement with Corollary 5.1, cannot be used to infer that $\boldsymbol{\mathbf{\psi}}$ has a finite Koopman mode expansion.
@@ 5.3.1 Van der Pol oscillator
Checking Corollary 5.1 for this system generates Fig. 2, where taking at-least $15$ time delays seems to produce a consistent step-like trend.
Figure 2: The system order, $r$, is unknown for the dictionary, $\boldsymbol{\mathbf{\psi}}$, that is used Eq. 21 to study the Van der Pol oscillator Eq. 20. Nonetheless, sufficient time delays lead to a step-like trend in Relative distance to DFT. By Corollary 5.1, the location of the jump might be indicative of $\boldsymbol{\mathbf{\psi}}$ possessing no more than 11 Koopman modes.
In particular, the jump in Relative distance to DFT occurs when $\theta$ increases from 10 to 11. Hence, $\boldsymbol{\mathbf{\psi}}$ might not possess more than 11 Koopman modes. Unfortunately, this speculation cannot be turned into a guarantee using Corollary 5.1.
@@ 5.3.2 Lid-driven cavity
Consider the lid-driven cavity for Reynolds numbers ($Re$) between $13\times 10^{3}$ and $30\times 10^{3}$. As $Re$ increases, the fluid flow transitions from periodic to chaotic, passing through quasi-periodic and mixed behavior [2]. We sample the continuous-time evolution [^15] Numerical simulations from [2]. of the lid-driven cavity at equi-spaced points in time to produce the discrete-time dynamics Eq. 1. We also choose an arbitrary linear functional of an associated stream-function as our observable. This sets the stage to study DMD-DFT equivalence, over a range of model orders $(\theta)$ and delay embedding dimension $(d)$.
Over the range of $Re$ considered, the qualitative change in dynamics is reflected in the relationship between Relative distance to DFT and $\theta$ (Fig. 3). As the Reynolds number increases, we see that the step-like trend has its jump smeared out over a wider range of $\theta$. The plateau begins at larger values of $\theta$ before disappearing altogether in the last scenario $(Re=30\times 10^{3})$. Hence, our observable may have a finite number of Koopman modes in the first three cases $(Re=13\times 10^{3},~{}16\times 10^{3}~{}\textrm{and}~{}20\times 10^{3})$.
However, for $Re=30\times 10^{3}$, we can go further and certify that our observable, $\boldsymbol{\mathbf{\psi}}$, possesses at least 25 Koopman modes. To get this guarantee, we first assume that $\boldsymbol{\mathbf{\psi}}$ has utmost 24 Koopman modes or, equivalently, that it lies in the span of utmost $24$ KEFs. In other words, we assume $r_{\rm max}=24$. The concomitant requirement of at least 23 delays is met by the study corresponding to the dark blue boxes in panel (d) of Fig. 3. Hence, by Corollary 5.1, we would expect it to spike, at the latest, by $\theta=25$. In contrast to studies with a lower number of time delays, the trend shown by the dark blue boxes lacks a jump. Hence, $\boldsymbol{\mathbf{\psi}}$ cannot have fewer than 25 Koopman modes.
@@
Figure 3: As $Re$ increases, the lid-driven cavity flow grows in complexity. This correlation is reflected in the above box plots of Relative distance to DFT. Panels (a)-(c) display jumps that occur at larger values of $\theta$. Furthermore, such jumps are eventually robust to the number of delays. According to Corollary 5.1, such a trend may be the result of an increase in the number of Koopman modes that comprise the observable $\boldsymbol{\mathbf{\psi}}$. In contrast, the final plot (d) is conspicuous in its lack of a discontinuity. By Corollary 5.1, we can infer that $\boldsymbol{\mathbf{\psi}}$ possesses at-least 25 Koopman modes. This observation agrees with the fact that the Koopman operator does not possess any eigenfunctions when the underlying dynamics is chaotic.
@@ 6 Conclusions and Future Work
For future work, we note that increasing the number of time delays, far beyond the prescription of Proposition 4.10, distributes the departure from DMD-DFT equivalence over two jumps. The theory developed in this work remains confined to predicting the first jump. Understanding the second jump along with the robustness of the overall transition to conditioning will improve the reliability of DMD-DFT equivalence in inferring the order of the underlying dynamics.
@@ Acknowledgments
Appendix A Unpacking the structure induced by Definitions 3.8 and 3.10
@@
    The dictionary $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ KEFs, $\{\,\phi_{i}\,\}_{i=1}^{r}$, with distinct eigenvalues, $\{\,\lambda_{i}\,\}_{i=1}^{r}$ (Definition 3.8).
@@
    The initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative (Definition 3.10).
In this setting, the Koopman mode expansion of $\boldsymbol{\mathbf{\psi}}$ (Definition 3.3) manifests as a specific factorization of the time series $\mathbf{Z}$ (Section A.1). Furthermore, one of these factors, the matrix of Koopman modes, provides an alternative characterization of the Koopman invariance (Definition 4.7) of $\boldsymbol{\mathbf{\psi}}$ (Section A.2).
A.1 The Koopman mode factorization of $\mathbf{Z}$
Lemma A.1.
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ KEFs, $\{\,\phi_{i}\,\}_{i=1}^{r}$, with distinct eigenvalues, $\{\,\lambda_{i}\,\}_{i=1}^{r}$.
Then, the row-space of $\mathbf{Z}$ is contained in the row space of the Vandermonde matrix,

### v6

@@ 5.1.2 Quantifying a theoretical loophole with ensemble studies
To paraphrase, $\mathbf{Z}$ corresponds to sampling $\boldsymbol{\mathbf{\psi}}$ on the first $\theta+d+1$ elements of the trajectory. Subsequently, we form $\mathbf{Z}_{\rm d-delayed}$ and compute Relative distance to DFT. This calculation is, then, iterated over every pair $(\theta,d)$ in the parameter regime of interest. Finally, we also repeat this parametric sweep, at a higher level, over many trajectories [^10] We also switch the dictionary for each trajectory for an added level of randomness.. Consequently, we can study the statistics of Relative distance to DFT for every pair of model-order ($\theta$) and delay-embedding dimension $(d)$ being tested. This is facilitated by box-whisker plots of this indicator with respect to model-order.
We test the major results on a variety of dynamical systems of increasing complexity. We primarily work with linear-time invariant (LTI) systems since their Koopman spectra are easily determined and the conditions required in Corollary 5.3 can be met. We also validate our theory using the Van der Pol oscillator and the lid-driven cavity from [2]. Here, in contrast to the LTI systems, we have no guarantee of meeting the requisites in Corollary 5.3 nor do we possess complete knowledge of the associated Koopman spectra. Nonetheless, we can still use the contrapositive of Corollary 5.3 to get a lower bound on the complexity of the underlying Koopman mode expansion.
@@
Eigenvalues $|\lambda_{i}|=1,~{}\forall\,i$ $|\lambda_{i}|=1\neq\lambda_{i},~{}\forall\,i$ $|\lambda_{i}|=1\neq\lambda_{i},~{}\forall\,i~{}\in~{}\{1,2,3,4\}$
$\lambda_{1}=1$ $|\lambda_{5}|,~{}|\lambda_{6}|,~{}|\lambda_{7}|~{}\in~{}[0.8,~{}0.9]$
@@
For the unitary systems $\textrm{LTI}_{1a}$ and $\textrm{LTI}_{1b}$, the expected trend is clearly visible in sub-figures (a) and (b) of Fig. 1. When $\theta$ is greater than or equal to $7~{}(8)$ for $\textrm{LTI}_{1a}~{}(\textrm{LTI}_{1b})$, the values of Relative distance to DFT are many orders of magnitude larger than those seen for smaller $\theta$. Moreover, the shadowing of this feature by the pertinent inter-quartile ranges (IQRs), which are denoted by boxes colored according to the number of delays taken, suggests that the observation is independent of the choice of initial condition.
@@
In contrast, the computations in $\textrm{LTI}_{3}$ exhibit a prominent dependence on the delay embedding dimension. Recall that Corollary 5.3 predicts the same trend for $\textrm{LTI}_{3}$ and $\textrm{LTI}_{1b}$ when at-least $6$ time delays are taken. In sub-figure (c) of Fig. 1, we observe that this is indeed true when only $6$ time delays are taken. However, when we increase the number of time delays to $24$, the spike in Relative distance to DFT at $\theta=8$ is mollified to the extent that it invites skepticism on the theoretical predictions.
This highlights the computational inadequacies of the forecasts given by Corollary 5.3. For further clarity, we need only formalize the criterion that we implicitly used to ascertain the validity of Corollary 5.3 in $\textrm{LTI}_{1a}$ and $\textrm{LTI}_{1b}$. Checking the predictions of Corollary 5.3 necessitates deciding what it means for Relative distance to DFT  to be “zero”. Specifically, we must select an $\epsilon\in(0,1]$ such that values of Relative distance to DFT  smaller than $\epsilon$ can be considered “zero”. For the unitary systems $\textrm{LTI}_{1a}$ and $\textrm{LTI}_{1b}$, we can choose any number between $10^{-14}$ and $10^{-10}$ as $\epsilon$, regardless of the delay embedding dimension $d$. However, in $\textrm{LTI}_{3}$, we can only pick $\epsilon$ in a narrow band around $10^{-10}$. Hence, the validity of Corollary 5.3 appears “forced” in $\textrm{LTI}_{3}$ when compared to $\textrm{LTI}_{1a}$ and $\textrm{LTI}_{1b}$.
@@
Figure 1: For the linear time-invariant (LTI) systems described by Eq. 31, box plots of Relative distance to DFT reveal its dependence on the model order $(\theta)$ and the number of time delays (Colour-coded). The magenta line indicates the system order $r$, which is $7$ for all three examples. When a minimum of 6 time delays are taken and the model order is at least 7 (8) for $\textrm{LTI}_{1a}(\textrm{LTI}_{1b})$, Relative distance to DFT  is larger than at lower model orders, as predicted by Corollary 5.3. The pertinent prediction for $\textrm{LTI}_{3}$ large values of Relative distance to DFT for $\theta\geq 8$ and $d\geq 6$ also appears to hold for $d=6$ i.e., when only the minimal required time delays are taken. However, increasing the delay embedding dimension to $24$ significantly diminishes the spike in Relative distance to DFT  at $\theta=8$, to the extent that it could be overlooked.
Remark 5.5.
Figure 1 shows secondary spikes in Relative distance to DFT for all three systems when an excess of time delays are taken. In particular, said spikes, which are present at $\theta=14,16~{}\textrm{and}~{}21$ for $\textrm{LTI}_{1a},\textrm{LTI}_{1b}~{}\textrm{and}~{}\textrm{LTI}_{3}$ respectively, are the result of $\left\lVert(\mathbf{X}_{\rm d-delayed})_{\rm ms}-(\hat{\mathbf{X}}_{\rm d-delayed})_{\rm ms}\right\rVert$ becoming non-trivial. Hence, by Remark 5.4, Fig. 1 depicts only an upper bound on Relative distance to DFT after the secondary spikes. For further discussion, see Appendix F.
@@
When we do not know if our dictionary $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs or if our initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative, we can still use the contrapositive of Corollary 5.3 to get a lower bound on the number of Koopman modes constituting our dictionary. First, we note that if the predictions of Corollary 5.3 are inconsistent with a numerical study that has taken sufficient time delays, then, we have either used a dictionary that is not non-redundantly spanned by a finite number of KEFs or our initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is not spectrally informative. In practice, the latter is not an issue because we can restrict our attention to only those KEFs whose zero level set does not contain $\boldsymbol{\mathbf{\upsilon}}_{1}$. This change of perspective is permissible because the reduced collection of KEFs is equally capable of generating the training set, $\mathbf{Z}$, through time-sequential observations. Since the condition of $\boldsymbol{\mathbf{\psi}}$ lying in the non-redundant span of $r$ distinct KEFs can be interpreted as requiring $\boldsymbol{\mathbf{\psi}}$ possess only $r$ Koopman modes, its negation, over the specific guesses of $r$ described below, can give a lower bound on the number of Koopman modes comprising $\boldsymbol{\mathbf{\psi}}$. Thus, when the delay embedding dimension, $d$, is large enough, numerical observations of Relative distance to DFT that are inconsistent with Corollary 5.3 can inform the number of Koopman modes necessary to represent $\boldsymbol{\mathbf{\psi}}$.
The contrapositive of Corollary 5.3 can be computationally leveraged by assuming an upper-bound on the number of Koopman modes, performing the ensemble experiment described in Section 5.1.2 and using errors, if any, in the predictions of Corollary 5.3 to negate the presumed upper-bound. In each forthcoming numerical experiment, we begin with the assumption that our observables are non-redundantly spanned by $r$ distinct KEFs. Although $r$ is unknown, we assume that it is finite and has a known upper bound of $r_{\rm max}$. This may be a theoretically informed cap or, if little is known about the system, simply the maximum model order that is palatable from a modeling perspective. Suppose we take at least $r_{\max}-1$ time delays i.e., $d\geq r_{\max}-1$. Then, by Corollary 5.3, Relative distance to DFT must behave like a step function with respect to the model order ($\theta$). It must be near zero for model orders less than $r+1$ ($r$ if 1 is a Koopman eigenvalue) and significantly higher for larger model orders. The absence of a step-like behavior for $\theta$ less than or equal to $r_{\max}+1$ would, then, mean that at-least $r_{\rm max}+1$ distinct KEFs are required to represent the span of our observables i.e., $\boldsymbol{\mathbf{\psi}}$ has at-least $r_{\max}+1$ Koopman modes. In contrast, the presence of a step-like behavior, in agreement with Corollary 5.3, cannot be used to infer that $\boldsymbol{\mathbf{\psi}}$ has a finite Koopman mode expansion.
@@ 5.3.1 Van der Pol oscillator
Checking Corollary 5.3 for this system generates Fig. 2, where taking at-least $12$ time delays seems to produce a consistent jump in Relative distance to DFT, when $\theta$ increases from 10 to 11.
Figure 2: The system order, $r$, is unknown for the dictionary that is used to study the Van der Pol oscillator Eq. 32. Nonetheless, sufficient time delays lead to a step-like trend in Relative distance to DFT. By Corollary 5.3, the location of the jump might be indicative of $\boldsymbol{\mathbf{\psi}}$ possessing no more than 11 Koopman modes.
Hence, $\boldsymbol{\mathbf{\psi}}$ might not possess more than 11 Koopman modes. Unfortunately, this speculation cannot be turned into a guarantee using Corollary 5.3.
@@ 5.3.2 Lid-driven cavity
Consider the lid-driven cavity for Reynolds numbers ($Re$) between $13\times 10^{3}$ and $30\times 10^{3}$. As $Re$ increases, the fluid flow transitions from periodic to chaotic, passing through quasi-periodic and mixed behavior [2]. We sample the continuous-time evolution [^12] Numerical simulations from [2]. of the lid-driven cavity at equi-spaced points in time to produce the discrete-time dynamics Eq. 1. We also choose an arbitrary linear functional of an associated stream-function as our observable. This sets the stage to study DMD-DFT equivalence, over a range of model orders $(\theta)$ and delay embedding dimensions $(d)$.
Over the range of $Re$ considered, the qualitative change in dynamics is reflected in the relationship between (the estimates [^13] See Appendix F of) Relative distance to DFT and $\theta$ (Fig. 3). As the Reynolds number increases, we see that the step-like trend associated with $d=25$ has its plateau beginning at larger values of $\theta$ before disappearing altogether in the last scenario $(Re=30\times 10^{3})$. Hence, our observable may have a finite number of Koopman modes in the first three cases $(Re=13\times 10^{3},~{}16\times 10^{3}~{}\textrm{and}~{}20\times 10^{3})$.
However, for $Re=30\times 10^{3}$, we can go further and certify that our observable, $\boldsymbol{\mathbf{\psi}}$, possesses at least 26 Koopman modes. To get this guarantee, we first assume that $\boldsymbol{\mathbf{\psi}}$ has utmost 25 Koopman modes or, equivalently, that it lies in the non-redundant span of utmost $25$ distinct KEFs. In other words, we assume $r_{\rm max}=25$. The concomitant requirement of at least 24 time delays is met by the study corresponding to the dark blue boxes in panel (d) of Fig. 3. Hence, by Corollary 5.3, we would expect it to spike, at the latest, by $\theta=26$. In contrast to the studies with a lower number of time delays ($d~{}=~{}3~{}\textrm{and}~{}12$), the trend shown by the dark blue boxes lacks a jump. Hence, $\boldsymbol{\mathbf{\psi}}$ cannot have fewer than 26 Koopman modes.
@@
Figure 3: As $Re$ increases, the lid-driven cavity flow grows in complexity. This correlation is reflected in the above box plots of (estimated) Relative distance to DFT. Panels (a)-(c) display jumps that occur at larger values of $\theta$. In contrast, the final plot (d) is conspicuous in its lack of a discontinuity. By Corollary 5.3, we can infer that $\boldsymbol{\mathbf{\psi}}$ possesses at-least 26 Koopman modes. This observation agrees with the fact that the Koopman operator does not possess any eigenfunctions when the underlying dynamics is chaotic.
@@ 6 Conclusions and Future Work
For future work, we note that increasing the number of time delays, far beyond the prescription of Proposition 4.10, can acutely blur the numerical distinction between DMD-DFT equivalence and non-equivalence. The theory developed in this work is oblivious to the choice of delay embedding dimension, beyond requiring a minimum value. Understanding this distortion produced by large delay embeddings will improve the reliability of DMD-DFT equivalence in inferring the order of the underlying dynamics.
@@ Acknowledgments
Appendix A Proof of Theorem 3.7
For any choice of $\boldsymbol{\mathbf{\upsilon}}_{\rm test}$, we can establish Eq. 18 by induction on $j$.
Step 1
When $j=1$, the identity Eq. 18 trivially holds.
Step 2
Suppose Eq. 18 holds for $j\leq o$.
Step 3
We need to show that Eq. 18 also holds for $j=o+1$ i.e.,
$\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1}~{}=~{}\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{1}.$
To this end, we will split the analysis into two cases:
Step 3.1: $o+1\leq n$
Sequential application of Eq. 16, Eq. 17 and Eq. 16 yields the desired conclusion.
$\mathbf{E}^{H}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1}}~{}=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}(\boldsymbol{\mathbf{z}}_{\rm test})_{o+1}}~{}=~{}\mathbf{E}^{H}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\boldsymbol{\mathbf{z}}_{\rm test})_{1}}~{}=~{}\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{1}.$
Step 3.2: $o+1>n$
Unpacking $(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1}$ with Eq. 16, we get:
$\mathbf{E}^{H}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1}}{}$ $=~{}\mathbf{E}^{H}\left(\boldsymbol{\mathbf{\mu}}_{\rm test}~{}+~{}\begin{bmatrix}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1-n+k}-\boldsymbol{\mathbf{\mu}}_{\rm test}\end{bmatrix}_{k=0}^{n-1}~{}~{}\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]\right)$
$=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}\boldsymbol{\mathbf{\mu}}_{\rm test}}~{}+~{}\begin{bmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1-n+k}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}\boldsymbol{\mathbf{\mu}}_{\rm test}}\end{bmatrix}_{k=0}^{n-1}~{}~{}\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}].$
The expression $\mathbf{E}^{H}\boldsymbol{\mathbf{\mu}}_{\rm test}$ can be simplified by applying Eq. 15 followed by Eq. 17:
$\mathbf{E}^{H}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{\mu}}_{\rm test}}~{}=~{}\frac{1}{n+1}\sum_{j=1}^{n+1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}(\boldsymbol{\mathbf{z}}_{\rm test})_{j}}~{}=~{}\frac{1}{n+1}\sum_{j=1}^{n+1}\mathbf{E}^{H}(\boldsymbol{\mathbf{z}}_{\rm test})_{1}~{}=~{}\mathbf{E}^{H}(\boldsymbol{\mathbf{z}}_{\rm test})_{1}.$
Similarly, the term $\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1-n+k}$ can be reduced by noting that $k\in\{0,\dots,n-1\}$. Since this means $o+1-n+k\leq o$, we can apply Step 2 of the induction followed by Eq. 16 to get:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1-n+k}}~{}=~{}\mathbf{E}^{H}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{1}}~{}=~{}\mathbf{E}^{H}(\boldsymbol{\mathbf{z}}_{\rm test})_{1}.$
Combining the above reductions and, then, using Eq. 16 yields the desired identity:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}\boldsymbol{\mathbf{\mu}}_{\rm test}}~{}+~{}\begin{bmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{o+1-n+k}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{E}^{H}\boldsymbol{\mathbf{\mu}}_{\rm test}}\end{bmatrix}_{k=0}^{n-1}~{}~{}\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]$
[…]
Appendix B Unpacking the structure induced by recurrent assumptions on the dictionary and the training trajectory
@@
    The dictionary $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs, $\{\,\phi_{i}\,\}_{i=1}^{r}$, with distinct eigenvalues, $\{\,\lambda_{i}\,\}_{i=1}^{r}$.
@@
    The initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative.
In this setting, the Koopman mode expansion of $\boldsymbol{\mathbf{\psi}}$ manifests as a specific factorization of the time series $\mathbf{Z}$ (Section B.1). Furthermore, one of these factors, the matrix of Koopman modes, provides an alternative characterization of Koopman invariant dictionaries (Section B.2).
B.1 The Koopman mode factorization of $\mathbf{Z}$
Lemma B.1.
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs, $\{\,\phi_{i}\,\}_{i=1}^{r}$, with distinct eigenvalues, $\{\,\lambda_{i}\,\}_{i=1}^{r}$.
Then, the row space of $\mathbf{Z}$ is contained in the row space of the Vandermonde matrix,